OpenStudy (anonymous):

[c++] .........Write a program that reads in the length and width of a rectangular yard (in meters) and the length and width of a rectangular house (in meters) placed in the yard. Your program should compute the time (in minutes) required to cut the lawn around the house. Assume the mowing rate in square meters per minutes is entered as a data input

OpenStudy (anonymous):

program conisits from input output an dprocess yours areall defined declare variables for length width mowing rate you find the area = l * w time = mowing rate * area thats all print time

OpenStudy (anonymous):

#includ<iostream.h> void main{ float l,w,area, mowrate; cout<<" Enter values for length & width"; cin>>l>>w; area=l * w; float time= mowrate * area; cout<<" The time taken is "<<time; }

OpenStudy (anonymous):

Make a correction the pre-processor directive is #include and the take values from the console for mowrate also . #include<iostream.h> void main() { float l,w,area, mowrate; cout<<" Enter values for length & width"; cin>>l>>w; area=l * w; cout<<" Enter rate of mowing in minutes "; cin>> mowrate; float time= mowrate * area; cout<<" The time taken is "<<time; }